# Lily Erickson EmerGen LLC Minneapolis, MN

### v2

@@ 6 Conclusion
Fortunately, based on this research alone as an indicator of where things currently stand, it seems unlikely that a modern machine learning algorithm will not be able to intuitively decipher the Elliptic Curve Discrete Logarithm Problem. Feeding a model cryptographically random information is an amazingly effective way to interfere with these kinds of attempts at backdooring modern curves and ciphers. Still, it’s likely only a matter of time before algorithms like these catch up to us.
@@
If we wanted to go completely overkill and use a 405B parameter model, we’d expect to invest an equivalent amount of effort to that required to break 3,000,000,000 (3 billion) keypairs the traditional way.
Now let’s conclude by asking ourself how much of the curve an attacker would actually need to invalidate in order to wreck widespread havoc over a population? 50%? 25%? I’m willing to bet that it’s less than 1%.
@@
Realistically, classical encryption needs orders of magnitude more complexity to survive. Simply doubling the key-length is no longer a sufficient deterrent.
Long-term, we need solutions in which there is not a perfect mapping between public and private key.
And with the groundwork laid, I’m willing to call the shot right here and now: classical 256-bit cryptography is dead. It’s only a matter of time before someone puts a super-computer on this problem and then leaks the model weights.
If you work in cyber-security and are not currently worried about AI, you should be. And if don’t, then you’d better be using a VPN these days.

### v3

@@ 5.1 The Birthday Paradox
In order to seriously discuss this problem, we are forced to dissect the above equation a bit. We propose two basic formulas for discussing the problem.
[…]
@@ 6 Conclusion
Fortunately, based on this research alone as an indicator of where things currently stand, it seems unlikely that a modern machine learning algorithm will not be able to intuitively decipher the Elliptic Curve Discrete Logarithm Problem. Feeding a model cryptographically random information is an amazingly effective way to interfere with these kinds of attempts at to backdoor modern curves and ciphers. Still, it’s likely only a matter of time before algorithms like these catch up to us.
@@
For more practical math, let’s assume that an attacker wants to compromise at least one of a generic S&P500 company’s credentials while having access only to their public keys in transit. To do this, they train a model with similar scale and scope to modern open-source AI on the marketplace right now, inline with our Llama 3 example.
Our model can consume significantly more tokens than traditional transformers because it uses a smaller embedding space, more efficient algorithm, and avoids the massive sequence length of language modeling. It also converges much faster than a language model, requiring fewer training epochs to memorize its data.
Additionally, there is no added burden of cleaning and annotating the data. Do not constrain yourself to the "one hundred trillion tokens" count that you might hear thrown around in the LLM world. The following example is akin to what an existing large research laboratory could create within a few months time.
Table 7: Practical Breakpoints
Parameter Value
Large Company’s Yearly Authentication Requests (Population) $200,000*365*500=7.3*10^{7}$
Has Memorized 1 Quantillion Keypairs (LLM-size) $10^{18}$
Current odds of solving a victim’s private key 0.00024%
If we want to assume a sufficiently motivated threat actor or nation-state surveillance organization, pushing the keypair count up to 1 sextillion ($10^{21}$) causes our odds of leaking an employee’s access credentials to skyrocket up at 0.14%.
I would also like to briefly touch on the fact that these are permanent solves of existing asymmetric encryption algorithms. Every year that an algorithm is commonly in use, there will be ever-evolving AI dedicated to solving them. Any sufficiently solved asymmetric encryption algorithm is then morphed into a symmetric key. As such, developers should begin to think about systems in which exposing a public key is a matter of last resort.
In the short term, using novel curves or having orders of magnitude more key complexity is a sufficient deterrent to prevent problems.
Long-term, we need solutions in which there is not a perfect mapping between a public and private key.
We are at the frontier of modern AI, and it has immense potential for both good and evil. It is my sincerest hope that everybody is doing everything in their power to keep themselves and their families safe. If you don’t use a modern VPN, now is a great time to start.